I 'm stuck on this part 19 3/4 - 9 9/10 =
Answer by stanbon(75887)

19 3/4 - 9 9/10 =
--------
(4*19+3)/4 - (10*9+9)/10
--------
79/4 - 99/10
-------------
= (10*79-4*99)/(4*10)
-------
= (790-396)/40
-------
= 9.85
============
Cheers,
Stan H.
============